OpenStudy (giovxnnx):

A DELL computer hardware has a yearly 17% rate of failure. Answer the following question. Round the answer up to six decimals; leave them in decimal form. What is the probability that, if you buy four (4) computers, at least one (1) computer will not fail?!

OpenStudy (giovxnnx):

Here is a picture of the question.

OpenStudy (dumbcow):

Let X be num of computers that will not fail. P(X >= 1) = 1 - P(X = 0) P(X=0) is probability that ALL 4 computers fail P(X>=1) = 1 - (.17)^4 = 0.99916 = 99.916%